### v2

# ARM-FM: Automated Reward Machines via Foundation Models for Compositional Reinforcement Learning

Abstract
Reinforcement learning (RL) algorithms are highly sensitive to reward function specification, which remains a central challenge limiting their broad applicability. We present ARM-FM: Automated Reward Machines via Foundation Models, a framework for automated, compositional reward design in RL that leverages the high-level reasoning capabilities of foundation models (FMs). Reward machines (RMs) – an automata-based formalism for reward specification – are used as the mechanism for RL objective specification, and are automatically constructed via the use of FMs. The structured formalism of RMs yields effective task decompositions, while the use of FMs enables objective specifications in natural language. Concretely, we (i) use FMs to automatically generate RMs from natural language specifications; (ii) associate language embeddings with each RM automata-state to enable generalization across tasks; and (iii) provide empirical evidence of ARM-FM’s effectiveness in a diverse suite of challenging environments, including evidence of zero-shot generalization.
1 Introduction
A central challenge in reinforcement learning (RL) is the design of effective reward functions for complex tasks. The shape of the reward influences the complexity of the problem at hand (Gupta et al., 2022); for instance, sparse rewards provide an insufficient learning signal, making it difficult for agents to improve (Devidze et al., 2022). Even hand-crafted dense rewards are susceptible to unintended loopholes or "reward hacking", where an agent exploits the specification without achieving the true objective (Fu et al., 2025). The unifying challenge is thus how to communicate complex objectives to an agent in a manner that provides structured, actionable guidance (Rani et al., 2025).
While Foundation Models (FMs) excel at interpreting and decomposing tasks from natural language, a critical gap exists in translating this abstract understanding into the concrete structured reward signals necessary for RL. Consequently, high-level plans generated by FMs often fail to ground effectively, leaving the agent without the granular feedback required for learning. To bridge this gap, we turn to Reward Machines (RMs), an automata-based formalism. By decomposing tasks into a finite automaton of sub-goals, RMs provide a compositional structure for both rewards and policies that is inherently more structured and verifiable than monolithic reward functions (Icarte et al., 2022). While theoretically principled, their practical application has been confined to task-specific applications due to the complexity of their manual, expert-driven design. We posit that the reasoning and code-generation capabilities of modern FMs are well-suited to automate the design and construction of RMs, thereby unlocking their potential to solve the broader challenge of communicating complex objectives in RL; the resulting RMs can thus translate abstract human intent into a concrete learning signal for solving complex tasks.
This work makes three primary contributions. First, we develop a novel framework for automatically generating complete task specifications directly from natural language using foundation models, introducing language-aligned reward machines (LARMs) which include the automaton structure, executable labeling functions, and natural language instructions for each subtask. Second, we introduce a method that leverages the language-aligned nature of the resulting automata to create a shared skill space, enabling effective experience reuse and policy transfer across related tasks. Finally, we provide extensive empirical validation demonstrating that our approach solves complex, long-horizon tasks across multiple domains that are generally intractable for standard RL methods. Specifically, our results show that the framework (i) dramatically improves sample-efficiency by converting sparse rewards into dense, structured learning signals; (ii) scales to a diverse set of environments, including grid worlds, complex 3D environments, and robotics with continuous control; and (iii) enables efficient multi-task training and zero-shot generalization.
Figure 1: An overview of our framework (left) and results in a complex sparse-reward environment (right). Reward Machine Generation (top-left): Given a high-level natural language prompt and a visual observation of the environment, a FM automatically generates the formal specification of the Reward Machine, the executable Python code for the labeling functions, and the natural language descriptions for each RM state. RL training (bottom-left): During the RL training loop, the labeling functions evaluate environment observations to update the Reward Machine’s state, which provides a dense reward signal $R_{t}^{\text{RM}}$. The RL agent’s policy receives the environment observation along with the embedding $\phi(\cdot)$ of the current RM state’s language description, making it aware of its active sub-goal. Empirical results (right): Results in a complex sparse-reward Minecraft-based resource-gathering task from Craftium (Malagón et al., 2024), where an RL agent is unable to make progress (top), while our agent, guided by an FM-generated LARM, learns to solve the task efficiently (bottom).
2 Automated Reward Machines via Foundation Models
Figure 2: ARM-FM leverage FMs to automatically construct RMs: using the UnlockToUnlock task description from MiniGrid (left), an RM is automatically constructed to solve the task (right).
We now present Automated Reward Machines via Foundation Models (ARM-FM), a framework for automated reward design in RL that leverages the reasoning capabilities of foundation models to automatically translate complex, natural-language task descriptions into structured task representations for RL training. Figure 1 illustrates an overview of ARM-FM, which comprises two major components: (i) the introduction of Language-Aligned RMs (LARMs), which are automatically constructed using FMs; and (ii) their integration into RL training by conditioning policies on language embeddings of RM states, enabling structured rewards, generalization, and skill reuse. Figure 2 shows a high-level task description, consisting of a natural language prompt and a visual observation (left), along with its corresponding RM (right). The RM is a finite-state automaton that guides an agent by providing incremental rewards for completing sub-goals, such as collecting keys and opening doors, on the way to the final objective. In the following section, we describe how our ARM-FM framework automates the creation of these reward machines directly from high-level task descriptions.
2.1 Language-Aligned Reward Machines
We assume the standard RL formalism, which defines an environment as a Markov Decision Processes (Puterman, 1994, ; MDPs) $\langle S,A,\mathcal{R},\mathcal{P}\rangle$, where $S$ is the set of MDP states, $A$ is the set of possible actions, $\mathcal{R}:S\times A\rightarrow\mathbb{R}$ is the MDP reward function, and $\mathcal{P}:S\times A\rightarrow\Delta(S)$ is a probabilistic MDP transition function. A Reward Machine (RM) is a finite-state automaton that encodes complex, temporally extended, and potentially non-Markovian RL tasks (Icarte et al., 2022). We formally define an RM by the tuple $\langle U,u_{I},\Sigma,\delta,R,F,\mathcal{L}\rangle$. Here, $U$ is the finite set of RM states; $u_{I}$ the initial state of the RM; $\Sigma$ is the finite set of symbols representing events that cause transitions in the RM; $\delta:U\times\Sigma\to U$ is the deterministic RM transition function; $R:U\times S\times A\times S\to\mathbb{R}$ is the RM reward function; $F\subseteq U$ is the set of final RM states; and $\mathcal{L}:S\times A\to\Sigma$ is the labeling function that connects MDP states $s\in S$ and actions $a\in A$ to the RM event symbols $\sigma\in\Sigma$. Intuitively, RMs are useful for describing tasks at an abstract level, especially when said tasks require multiple steps over long time horizons. Each RM state $u\in U$ can be thought of as representing a subtask, and the transitions $u^{\prime}=\delta(u,\sigma)$ denote progress to a new stage of the overall objective after a particular event $\sigma\in\Sigma$ occurs in the environment. The RM reward function $R(u,s,a,s^{\prime})$ assigns a reward based on the current RM state $u$ and the underlying MDP transition $(s,a,s^{\prime})$. Meanwhile, the set of final RM states $F$ defines the conditions under which the task described by the RM is complete. Finally, the labeling function $\mathcal{L}$ is required to connect the RM’s events, transitions, and rewards to states and actions from the underlying MDP.
Figure 3: A self-improvement loop where a generator and critic FMs iteratively refine LARMs, with optional human verification.
We define LARMs as RMs that are additionally equipped with natural-language instructions $l_{u}$ for each RM state $u$, and with an embedding function $\phi(\cdot)$ that maps such language instructions to an embedding vector $z_{u}=\phi(l_{u})\in\mathbb{R}^{d}$. We note that by equipping RM states with embedding vectors $z_{u}$ that encode language-based descriptions of the corresponding subtasks, we provide the first mechanism for constructing a semantically grounded skill space in RMs: policies conditioned on these embeddings can naturally share knowledge across related subtasks, enabling transfer, compositionality, and zero-shot generalization.
Figure 4: The three core components generated by our method for the UnlockToUnlock environment: (Left) the RM specification, (Center) the labeling functions that drive the state transitions, and (Right) the instructions and embeddings for each RM state.
We present a framework to automatically construct LARMs from language-and-image-based task descriptions by iteratively prompting an FM, as is illustrated in Figure 3. More specifically, to progressively refine the RM specification, we employ N rounds of self-improvement using paired generator and critic FMs (Tian et al., 2024). In practice, we find that FM-generated reward machines are both interpretable and easily modifiable, as they follow a natural language specification. Figure 4 illustrates an automatically-constructed LARM for the UnlockToUnlock task, including a text-based description of the RM (left), FM-generated labeling functions $\mathcal{L}$ (middle), and natural RM state instructions and embeddings $l_{u}$ (right). All RMs and labeling functions used in this work are shown in Appendix 6.6 and 6.7.
2.2 Reinforcement Learning with LARMs
The introduction of the LARM uses an augmented state space that is the cross-product of the MDP and RM states ($\mathcal{S}\times\mathcal{U}$), and a reward function that is the sum of the MDP and RM rewards; we will refer to this augmented MDP as $\mathcal{M}^{\prime}$, and it is illustrated in Figure 1 (Bottom). At timestep $t$, the agent selects actions conditioned on the environment state and the language embedding of the current LARM state: $\pi(s_{t},z_{u_{t}})$. This language-based policy conditioning is the central mechanism enabling generalization in our framework, creating a semantically grounded skill space where instructions like “pick up a blue key" and "pick up a red key" are naturally close in the embedding space, unlocking a pathway for broad experience reuse and efficient policy transfer.
During training, after the agent executes an action $a_{t}\sim\pi(s_{t},z_{u_{t}})$, the underlying MDP transitions to $s_{t+1}$ and returns a reward $R_{t}$. The labeling function $\mathcal{L}(s_{t+1},a_{t})$ determines if a symbolic event has occurred, which may induce a LARM transition $u_{t+1}=\delta(u_{t},\mathcal{L}(s_{t+1},a_{t}))$, as well as an additional reward $R^{\text{RM}}_{t}$. The sum of the MDP and RM rewards are then used for learning: $R^{\text{total}}_{t}=R_{t}+R^{\text{RM}}_{t}$. This complete training procedure, adapted for a DQN agent, is formalized in Appendix 6.3.
3 Empirical Results
We present a series of experiments designed to evaluate the effectiveness and scalability of our method: we test generalization and long-horizon planning in sparse-reward settings with the MiniGrid and BabyAI suites (Chevalier-Boisvert et al., 2023) (Section 3.1), we evaluate scalability with a resource-gathering task in a 3D, procedurally generated Minecraft world from Craftium (Malagón et al., 2024) (Section 3.2), and we demonstrate the applicability of our approach to create RMs that work in continuous control in challenging robotics tasks from Meta-World (McLean et al., 2025) (Section 3.3). Finally, we use XLand-MiniGrid (Nikulin et al., 2024) to evaluate the generalization capabilities of RL agents trained with LARMs (Section 3.4). Screenshots of these environments are shown in Figure 5 and environment-specific details in Appendix 6.2.
We report results averaged over 3 independent random seeds, with shaded regions and error bars indicating one standard deviation. Comprehensive details for all environments as well as additional results are presented in Appendix 6.2, details on the baselines used and additional ablations in 6.4, and hyperparameters in 6.8.
(a) MiniGrid
(b) Craftium
(c) Meta-World
(d) XLand-MiniGrid
Figure 5: An overview of the diverse environments used to evaluate ARM-FM. (a) MiniGrid environments test long-horizon planning in sparse-reward settings. (b) Craftium, a 3D world inspired by Minecraft, evaluates how ARM-FM scales with increasing task complexity. (c) Meta-World provides a benchmark for robotic manipulation with continuous control. (d) XLand-MiniGrid is used to evaluate multi-task learning and generalization.
3.1 Sparse Reward Tasks
We first evaluate our method on the MiniGrid suite of environments, which are challenging due to sparse rewards. For these experiments, we use DQN (Mnih et al., 2013) as the base agent and compare our method (DQN+RM) with a baseline with intrinsic motivation (DQN+ICM) (Pathak et al., 2017), an LLM-as-agent baseline from Paglieri et al. (2024), as well as an unmodified DQN. In Appendix 6.4.4 we include results comparing to a VLM-as-reward-model baseline proposed by Rocamonde et al. (2023). The LARMs used to train our DQN+RM agent are shown in Appendix 6.6.
Figure 6: Performance on MiniGrid-DoorKey environments of increasing size. The top row shows performance on a fixed map layout, while the bottom row shows performance on procedurally generated layouts. Our method consistently achieves higher rewards across all tasks.
Our agent successfully solves a suite of challenging exploration tasks where all baselines fail. We first present results on the DoorKey task in Figure 6, showing performance across increasing grid sizes. Our method consistently outperforms the baselines in all settings, including fixed maps (top row) and procedurally generated maps where the layout is randomized each episode (bottom row). We provide an additional analysis of our method’s success in Appendix 6.4.2.
To further test our agent, we select three significantly harder tasks that require longer planning horizons: UnlockToUnlock, BlockedUnlockPickup, and KeyCorridor. As demonstrated in Figure 7, our agent is the only one capable of solving all three tasks and achieving near-perfect reward, while all other baselines fail to make any progress.
Figure 7: Performance on complex, long-horizon MiniGrid tasks. Our method successfully solves all three, while baselines show no learning.
3.2 Scaling to Complex 3D Environments
We assess our method’s performance in a complex, procedurally generated 3D environment using Craftium (Malagón et al., 2024), which is a Minecraft-based resource-gathering task. Here, the agent’s goal is to mine a diamond by first navigating the world to gather the required wood, stone, and iron. The environment provides a sparse reward only upon collecting the final diamond. For this set of experiments we use PPO (Schulman et al., 2017) as the base agent.
As shown in Figure 1 (right), PPO augmented with our generated LARM consistently completes the entire task sequence, while the baseline PPO agent makes minimal progress. This result is particularly significant, as RMs often require manual, expert-driven design which can become challenging in complex, open-ended environments (Icarte et al., 2022). In contrast, we demonstrate the successful application of a RM that is not only automatically generated by a FM but is also highly effective in a complex 3D, procedurally generated environment. The LARM achieves this by effectively decomposing the high-level goal into progressive subtasks and providing crucial intermediate rewards. This experiment highlights our framework’s ability to handle increased action dimensionality and visual complexity, and it showcases the capability of FMs to leverage their knowledge to automate task decomposition.
3.3 Robotic Manipulation
Our framework can automate the complex task of reward engineering for robotic manipulation, providing dense supervision with a FM-generated LARM. We evaluate this capability in continuous control domains using the Meta-World benchmark (McLean et al., 2025), where designing dense reward functions typically requires extensive hand-engineering of low-level signals (e.g., joint angles). Our approach bypasses this difficulty entirely. The resulting reward machine offers richer learning signals than sparse rewards, enabling the agent to make more progress. As demonstrated in Figure 8, our method achieves higher success rates than learning from sparse rewards alone, using SAC (Haarnoja et al., 2018) as the base agent. Additional experiments on Meta-World are provided in Appendix 6.2.3.
Figure 8: Performance on Meta-World manipulation tasks. In most tasks, our method achieves high success rates compared to the sparse reward agent.
3.4 Generalization through language embeddings
Figure 9: Ablation study on the components of our framework. A single Rainbow agent is trained on an increasing number of simultaneous tasks. While the baseline fails to generalize, and ablations show partial success, only our full method, which uses both LARM rewards and state embeddings, maintains a high success rate as the number of tasks grows.
A key design choice in our framework is to condition the agent’s policy on the language embeddings of the current RM state, which contrasts with prior work that uses separate policies that do not permit knowledge sharing (Alsadat et al., 2025). In the following, we (i) ablate the roles of the LARM rewards and state embeddings in enabling an RL agent to learn robust, multi-task policies; and (ii) demonstrate how the compositional structure of LARMs leads to zero-shot generalization on unseen tasks.
Both structured rewards and language-based state conditioning are essential for learning a robust, multi-task policy. To disentangle the benefits of the LARM’s reward structure from the state embeddings, we conduct an ablation study in a multi-task setting. We train a single Rainbow DQN agent on an increasing number of simultaneous XLand-MiniGrid tasks and measure its average success rate. As shown in Figure 9, the baseline agent fails to generalize as the number of tasks increases. Providing the policy with only the state embeddings gives it a weak learning signal that degrades quickly. Conversely, providing only the LARM rewards enables multi-task learning, but the policy struggles as it is unaware of the active sub-goal. Our full method, which uses both the dense rewards from the LARM and the state embeddings to condition the policy, is robust and maintains high performance even when trained on 10 simultaneous tasks.
The compositional structure of LARMs enables zero-shot generalization to novel tasks composed of previously seen sub-goals. The ultimate test of our compositional approach is whether the trained policy, $\pi(a_{t}|s_{t},z_{u_{t}})$, can solve a novel task without any additional training. We design an experiment where $\pi$ is trained on a set of tasks, $\{\mathcal{T}_{A},\mathcal{T}_{B}\}$, each with an associated LARM, $\mathcal{R}_{A}$ and $\mathcal{R}_{B}$. During training, the policy learns skills corresponding to the union of all sub-goal embeddings, $\{z_{u}|u\in U_{A}\cup U_{B}\}$. At evaluation time, we introduce a new, unseen task, $\mathcal{T}_{C}$, with a novel LARM, $\mathcal{R}_{C}$, generated by the FM. Zero-shot success is possible if the set of sub-goals in the new task is composed of elements semantically familiar from training, i.e., if for any state $u^{\prime}\in U_{C}$, its embedding $z_{u^{\prime}}$ is close to an embedding seen during training. As illustrated in Figure 10, the agent successfully solves Task C. When the LARM for Task C transitions to a state $u^{\prime}_{t}$, the policy receives the input $(s_{t},z_{u^{\prime}_{t}})$. Because the embedding $z_{u^{\prime}_{t}}$ (e.g., for "Pick up a blue key, Position yourself to the right of the blue pyramid") is already located in a familiar region of the skill space, the policy can reuse the relevant learned behavior to make progress and solve the unseen composite task.
Figure 10: Demonstration of zero-shot generalization. An agent is trained on a set of tasks (A, B). At evaluation, it is given a new LARM for an unseen composite task (C). Because the sub-tasks in C (e.g., "Pick up a blue key") are semantically familiar from training, the agent can reuse learned skills to solve the novel task without any fine-tuning.
4 ARM-FM: In-depth analysis
We now conduct a fine-grained analysis of our method’s key components by evaluating (i) the quality of the LARMs generated by different FMs; and (ii) the semantic structure of the state embeddings.
(a) FM generation correctness.
(b) Semantic structure of state embeddings.
Figure 11: Analysis of FM-generated task components. (a) An LLM-as-judge evaluation across 1,000 tasks reveals a strong scaling trend, where larger foundation models more reliably generate correct RM structures and verifier code. (b) PCA visualization of thousands of state instruction embeddings shows a clear semantic structure, with start, middle, and end states forming distinct clusters.
Larger foundation models generate syntactically correct task structures with significantly higher reliability. To evaluate the FM’s generation capabilities, we sampled 1,000 diverse tasks from the XLand-MiniGrid environment (Nikulin et al., 2024) and prompted various open-source models to generate the corresponding reward machines, Python labeling functions, and natural language instructions. We compare models of different scales from the Qwen (Qwen, 2025), Gemma (Gemma, 2025), Llama (Dubey et al., 2024) and Mistral families. We employed an LLM-as-judge protocol to score the correctness of the generated artifacts (Gu et al., 2024), using Qwen3-30B-A3B-Instruct-2507 as the judge (a FM not used to generate the LARMs). Figure 11(a) shows a clear scaling trend: larger models like Qwen3-32B are significantly more capable of generating fully correct task specifications. Interestingly, some models exhibit different strengths; for instance, Mistral-Small is more adept at generating a valid RM structure than correct labeling code, highlighting the distinct reasoning and coding capabilities required.
The FM-generated state instructions produce a semantically coherent embedding space that clusters related sub-goals. Beyond syntactic correctness, the agent’s ability to generalize depends on the semantic quality of the state instruction embeddings. A well-structured embedding space should group semantically similar sub-tasks together, regardless of the overarching task. We analyze this by visualizing the embeddings of state instructions from the 1,000 generated XLand-MiniGrid tasks using PCA (Abdi and Williams, 2010), as shown in Figure 11(b). We used Qwen3-30B-A3B-Instruct-2507 to obtain the embeddings. The embeddings form distinct and meaningful clusters, with instructions corresponding to the start, middle, and end of a task occupying different regions of the space. Notably, instructions with similar meanings from different tasks cluster together, confirming that the FM produces a coherent representation. This underlying semantic structure is what enables a shared policy to treat related sub-tasks in a similar manner, forming the foundation for skill transfer.
5 Related Work
Reward Machines in Reinforcement Learning. RMs are a formal language representation of reward functions that expose the temporal and logical structure of tasks, thus enabling decomposition, transfer, and improved sample efficiency in learning (Icarte et al., 2018, 2022). For these reasons, RMs and related formal methods for task specification have been applied to address diverse challenges, from multiagent task decomposition (Neary et al., 2021; Smith et al., 2023) to robotic manipulation and task planning (Camacho et al., 2021; He et al., 2015; Cai et al., 2021). Recent work continues to broaden their applicability, by studying extensions that increase their expressivity (Varricchione et al., 2025), and by addressing uncertainty in symbol grounding and labeling functions (Li et al., 2024, 2025). While RMs can be difficult to design for non-experts, Toro Icarte et al. (2019) and Xu et al. (2020) propose methods that simultaneously learn RMs and RL policies, if the RM is unknown a priori. More closely related to our work, Alsadat et al. (2025) recently proposed to use language models to generate RMs for RL training. However, the above works treats RM states as isolated symbols, requiring carefully-defined mappings between states in different RMs to enable policy re-use. By contrast, ARM-FM not only generates RMs from natural language task descriptions, but also introduces a natural mechanism for connecting RM states by embedding their associated subtask descriptions in a shared latent space. Conditioning the policy on these language embeddings can thus enable knowledge transfer across similar subtasks, even when they occur in different RMs.
Foundation Models in Decision-Making. The emergence of FMs has inspired two main lines of work in sequential decision-making. The first uses FMs directly as autonomous agents (Paglieri et al., 2024). Approaches such as ReAct (Yao et al., 2023) and Voyager (Wang et al., 2023) employ large language models (LLMs) to perform reasoning, planning, and acting. While these systems demonstrate strong capabilities in complex domains, they heavily depend on environment abstractions (e.g., textual interfaces or code as actions) that bypass many of the low-level perception and control challenges central to RL.
A second line of research integrates FMs with RL training by using them to provide auxiliary signals such as high-level goals (Ahn et al., 2022) or reward feedback. For example, Motif (Klissarov et al., 2023) elicits trajectory-level preferences from FMs and distills them into a reward model. ONI (Zheng et al., 2024) aggregates asynchronous LLM feedback into a continuously updated reward function. Eureka (Ma et al., 2023) leverages evolutionary strategies to generate programmatic reward functions, which are then used to train downstream policies. ELLM uses pretrained LLMs to suggest plausibly useful goals and trains RL agents with goal-reaching rewards. These approaches illustrate the potential of injecting FM knowledge to shape RL objectives. However, the outputs are typically limited to an opaque reward model, rather than a structured, compositional representation of the task.
Our work differs in the structure of the FM–RL interface. We employ FMs to generate language-aligned RMs: structured, compositional, and interpretable representations of task reward functions. This formulation combines the expressivity of FMs with the explicit, modular decomposition, and human-in-the-loop refinement enabled by RMs, offering a principled path toward hierarchical and interpretable RL. Additionally, our method does not depend on specific environment abstractions (Wang et al., 2023) or the availability of a natural language interface (Klissarov et al., 2023).
Exploration. Sparse rewards present a fundamental exploration challenge in RL, as discovering rewarding states via random search is often intractable in high-dimensional environments (Sutton et al., 1998). Classic strategies based on state-visitation counts or optimism provide theoretical guarantees but scale poorly to complex state spaces where such counts are infeasible (Strehl and Littman, 2008; Auer et al., 2008; Bellemare et al., 2016). A dominant paradigm in deep RL is intrinsic motivation (Yuan et al., 2024), which augments sparse rewards with a dense, task-agnostic signal. These methods are typically either knowledge-based, rewarding an agent for improving its world model’s predictions (Pathak et al., 2017; Schmidhuber, 1991; Sekar et al., 2020), or novelty-based, rewarding visits to unfamiliar states (Bellemare et al., 2016; Burda et al., 2018; Ostrovski et al., 2017). However, their task-agnostic nature can lead to inefficient exploration misaligned with the true objective, a failure mode known as the “noisy TV” problem (Friston et al., 2012). To make exploration more directed, goal-oriented and hierarchical methods reframe the problem as reaching self-generated goals or learning a repertoire of skills (Andrychowicz et al., 2017; Eysenbach et al., 2018; Klissarov and Machado, 2023). While more structured, these typically rely on complex unsupervised objectives to discover useful goals, which can be unstable (Klissarov et al., 2025). In contrast, our work leverages explicit knowledge from natural language to construct a task-aware reward and state structure, avoiding the pitfalls of purely task-agnostic or unsupervised exploration.
6 Conclusion
In this work, we introduce Automated Reward Machines via Foundation Models (ARM-FM), a framework that bridges the critical gap between the semantic reasoning of foundation models and the low-level control of reinforcement learning agents. Our central contribution is a method for automatically generating Language-Aligned Reward Machines (LARMs) from natural language. We demonstrated that by conditioning a single policy on the embeddings of the LARM’s natural language state descriptions, we transform the reward machine from a static plan into a compositional library of reusable skills. Our experiments confirmed the effectiveness of this approach. We showed that ARM-FM solves a suite of long-horizon, sparse-reward tasks across diverse domains – from 2D grid worlds to a procedurally generated 3D crafting environment – that are intractable for strong RL baselines. Our analysis revealed that this performance is underpinned by a coherent semantic structure in the state embedding space and that both the structured rewards and the state embeddings are critical for robust multi-task learning. The ultimate validation of our compositional approach was the demonstration of zero-shot generalization to a novel, unseen task without any additional training.
Ultimately, this work establishes language-aligned reward machines as a powerful and versatile framework connecting foundation models, RL agents, and human operators. The modular, language-based structure allows FMs to generate accurate plans, agents to learn generalizable skills, and humans to easily inspect and refine the task specifications. While this paradigm is promising, one tradeoff of our approach is the human verification step during RM generation. On one hand, this step may be viewed as a feature – the language-based reward structures output by ARM-FM provide an interface for humans to interpret and refine task specifications. On the other hand, this step presupposes access to human verifiers. We note, however, that such verifiers are not strictly required, although they can improve output quality when available. Furthermore, future work may reduce or eliminate this dependence by exploiting the automaton-based structure of RMs to enable automated self-correction, for example, through formal verification. More broadly, we believe this work paves the way for a new class of RL agents that can translate high-level human intent and FM-generated plans into competent, generalizable, and interpretable behavior.